OpenStudy (anonymous):

Write 3/5 as a decimal and add it to 6.5. What is the sum? First one to answer gets a medal?

OpenStudy (anonymous):

thanks

myininaya (myininaya):

\[\frac{3}{5}+6.5=0.6+6.5=7.1\]

myininaya (myininaya):

Please @madbird2011 at least ask for an explanation. Please try to understand the material.